can i please have help in solving this linear function. f(x) has slope-2/3 and contains the point f(-3)=0. What is f(0)?
Answer by nerdybill(7384)

f(x) has slope-2/3 and contains the point f(-3)=0. What is f(0)?
.
slope intercept form of a line:
y = mx + b
to find 'b', plug in the given info from the problem:
m is -2/3
y is 0
x is -3
.
y = mx + b
0 = (-2/3)(-3) + b
0 = 2 + b
-2 = b
.
So, our equation of the line is:
f(x) = (-2/3)x - 2
.
To find f(0), plug it in:
f(x) = (-2/3)x - 2
f(0) = (-2/3)0 - 2
f(0) = -2 (notice, this is the same as 'b' (y-intercept) found earlier)